A piece of piping is the shape of a cylinder
The piping is rolling on a horizontal floor
Every time the piping completes 6 full rotations it travels a distance of 1 meter
What is the diameter of the Cylinder in mm

Answer by TimothyLamb(4379)

c = 2*pi*r
c = pi*d
d = c/pi
---
c = 1000/6 mm
---
d = (1000/6)/pi
d = (1000/6)/3.1415
d = 53.053212372 mm
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---